OpenStudy (anonymous):

Having a little trouble on a calc II problem. I'm trying to find the inverse of a function like this: f(x)=(2x^2+5x)^(1/2). I started out by finding f'(x), and then (f^1)'(x)=1/f'(x). Now with the slope of the inverse function, I integrated the reciprocal slope for...the wrong answer. Oops. Is this the right approach?

OpenStudy (isaiah.feynman):

Finding the inverse of a function is a basic calc 1 problem.

OpenStudy (anonymous):

Well, it ended up in my Calc II course.

OpenStudy (isaiah.feynman):

Procedure for finding inverse of a function Write f(x) as y Switch x and y Solve for y New function acquired is the inverse

OpenStudy (anonymous):

Yeah, I tried it that way (algebraically), to. Maybe I'm just making an algebra mistake.

OpenStudy (anonymous):

I meant "too".

OpenStudy (isaiah.feynman):

That's the only way I have seen it in textbooks.. Anyways try again.

OpenStudy (anonymous):

The problem is, on a function like this, you can't neatly isolate y.

OpenStudy (isaiah.feynman):

I kinda saw that. I will try a back of the envelope calculation.

OpenStudy (anonymous):

I think you can get there by switching x and y without solving. Differentiating both sides gives dx/dy, which integrated ought to give the inverse...I think.

OpenStudy (anonymous):

can someone please help me

OpenStudy (anonymous):

Isaiah, I think that's it. I can see a u substitution opening up. Daniella14, I'd recommend starting a new question thread.

OpenStudy (anonymous):

can i give you the question here ?

OpenStudy (anonymous):

@eighthourlunch

OpenStudy (anonymous):

I'd rather you didn't. More people will see it if you add it to the main board, and it won't mess up the one I'm working on.

OpenStudy (anonymous):

@eighthourlunch have you tried completing the square first?

OpenStudy (anonymous):

Ah, I haven't. I'l try that next, thanks. :)

OpenStudy (freckles):

You can even write as y = ax^2+bx+c as 0= ax^2+bx+(c-y) and then use quadratic formula to solve for x. Then replace x with f^-1(x) and replace y with x

OpenStudy (freckles):

However you need to take the restrictions of the domain and range of the original function into consideration for choosing the right inverse function

OpenStudy (anonymous):

SithsAndGiggles, that was it, thank you! I'd have tried it sooner, but my wife and I were just leaving for our 17th anniversary dinner. I thought it best to prioritize. :) Love your username, btw. Thanks to everyone else, as well.

OpenStudy (anonymous):

You're most welcome!